OpenStudy (anonymous):

Two perpendicular chords both 5cm from the circle divides the circle into four parts. Compute the area of the sector containing the smallest part if the radius of the circle is 13cm

OpenStudy (ranga):



OpenStudy (ranga):

Area of sector CBE = Area of sector OCB - Area of triangle OCB + Area of triangle CEB

OpenStudy (ranga):

FB = sqrt(OB^2 - OF^2) = sqrt(13^2 - 5^2) = sqrt(144) = 12 FE = 5 (given) EB = FB - FE = 12 - 5 = 7 Similarly, EC = 7

OpenStudy (ranga):

CB = sqrt(EB^2 + EC^2) = sqrt(7^2 + 7^2) = sqrt(2*7^2) = 7sqrt(2) Angle COB can be found from the law of cosines: CB^2 = OC^2 + OB^2 - 2(OC)(OB)cosCOB 49*2 = 13^2 + 13^2 - 2*13^2*cos(COB) Find angle COB Area of sector COB will be: Area of circle * angle COB / 360 Area of triangle COB = 1/2 * 13 * (13 * sin(angle COB)) Area of triangle CEB = 1/2 * EB * EC = 1/2 * 7 * 7 Put them all in the formula for the the area of sector CBE

OpenStudy (anonymous):

Thank you! I finally got it..

OpenStudy (ranga):

you are welcome.